Circular No. 7557

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001D IN IC 728 M. Modjaz and W. D. Li, University of California at Berkeley, on

behalf of LOTOSS (cf. IAUC 7514), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Jan. 5.5 (mag ∼ 17.6) and 6.4 UT (mag ∼ 17.5). The new object is located at α = 11^{h}44^{m}50^{s}.97, δ = −1^{o}36′18′′.4 (equinox 2000.0), which is 8′′.0 east and 12′′.4 south of the nucleus of IC 728. A KAIT image of the same field taken on 2000 Dec. 28.5 showed nothing at the position of SN 2001D (limiting mag ∼ 18.5).

SUPERNOVA 2001E IN NGC 3905 Modjaz and Li also report the discovery with KAIT of another appar-

ent supernova on unfiltered images on Jan. 5.5 (mag ∼ 17.6) and 6.4 UT (mag ∼ 17.2). SN 2001E is located at α = 11^{h}49^{m}05^{s}.00, δ = −9^{o}44′10′′.9 (equinox 2000.0), which is 0′′.6 east and 23′′.2 south of the nucleus of NGC 3905. A KAIT image of the same field taken on 2000 Dec. 30.5 showed nothing at the position of the new object (limiting mag ∼ 19.0).

SUPERNOVA 2000fo IN PGC 70148 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2000fo, obtained by M. Calkins on 2000 Dec. 23.08 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova at maximum light. The spectrum exhibits a possibly reddened continuum, with well-defined Si II, S II, Fe II, Mg II, and Ca II features. Adopting the NASA/IPAC Extragalactic Database recession velocity of 7152 km/s for the host galaxy, the expansion velocity is 11 800 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 3 ± 2 days before maximum light.

POSSIBLE NOVA IN PUPPIS CCD photometry of the new variable star noted on IAUC 7552 and

7556 (comparison stars from Tycho catalogue): 2000 Dec. 31.001 UT, V = 9.24 ± 0.05, B − V = +0.44 ± 0.07 (D. Hanžl, Brno, Czech Republic); 31.96, V = 9.14 ± 0.06, B − V = +0.48 ± 0.06 (Hanžl); 2001 Jan. 1.020, V = 9.13 (M. Lehký, Hradec Králové, Czech Republic).

2001 January 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green